Circular No. 5887

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993ad IN IC 1501 C. Pollas, Observatoire de la Côte d’Azur, reports his discovery of a

supernova of mag B = 18 on two films taken Nov. 7.87 and 8.88 UT. The object is located at α = 23^{h}32^{m}06^{s}.75, δ = −3^{o}26′08′′.5 (equinox 1950.0), which is 13′′ east and 22′′.7 south of the nucleus of the spiral host galaxy. Nothing was found at this position to mag 20 on films taken Oct. 20.88 and to mag 21 on Aug. 19.09. A nearby star is at end figures 12^{s}.89, 23′52′′.8. E. Cappellaro and M. Della Valle, European Southern Observatory, report: “A flux-calibrated CCD spectrogram (range 380-680 nm, resolution ∼ 1 nm), obtained on Nov. 10.1 UT with the 3.6-m reflector (+ EFOSC) at La Silla, shows this object to be a type-II supernova at a very early stage. Preliminary analysis reveals emission lines of the Balmer series from Hα to Hϵ superimposed on a very blue continuum. Other emission lines are measured at 462, 476 (likely He II 468.6), and 590 nm. All features are narrow (FWHM ∼ 1500 km/s). The spectrum bears strong similarity to that of SN 1983K ten days before maximum (Niemela et al. 1985, Ap.J. 289, 52). Provisional photometry gives B = 16.3, B–V = −0.3, V–R = 0.0. The Hα emission of the galaxy gives a redshift of z ∼ 0.018.”

CRAB NEBULA O. C. De Jager and H. I. Nel, Potchefstroom University, South Africa;

A. K. Harding, P. Sreekumar, and D. J. Thompson, Goddard Space Flight Center, NASA; and P. F. Michelson and P. L. Nolan, Stanford University, report: “EGRET γ-ray observations of the Crab unpulsed component yielded the following flux levels (units of 10^{−6} cm^{−2} s^{−1}) for the energy range 0.07-1 GeV: 1991 Apr. 23-May 30, 1.18 ± 0.11; 1992 June 11-1993 May 24, 0.58 ± 0.14. The significance of the reduction in flux is 3.4σ. The energy spectrum of the range 1-30 GeV is much flatter, and no reduction in flux is seen for this range. The variability below 1 GeV (as predicted by De Jager and Harding 1992, Ap.J. 396, 161) may be associated with unstable electron acceleration at the pulsar wind shock at a distance of 8′′ from the pulsar (the ‘Crab wisps’). EGRET observations of the Crab region continue between 1993 Dec. 1 and 13, and we urge optical observations (e.g., CCD) of the central Crab region. The measurement of temporal broadening from radio pulsar scintillation (cf. Gwinn et al. 1993, Ap.J. 410, 673) is also recommended.”

1993 November 10 Daniel W. E. Green